OpenStudy (osuxbucks15):

Please Help...thermochemical equation How many ML of water at 15 degrees C must be added to 85 ML of water at 65 degrees C to obtain a final temperature of 29 degrees C. ( Assume no heat was lost to surroundings)

OpenStudy (osuxbucks15):

@Mertsj can you help?

OpenStudy (mertsj):

There is some formula ( I don't remember it) that deals with this. Based on the question, you should have the formula in close proximity.

OpenStudy (osuxbucks15):

is it q= m x C x delta T?

OpenStudy (anonymous):

yep thats the equation

OpenStudy (anonymous):

notice that you need mass and the measurements in the problem are using volume/ so we have to use "density" to do the conversion there

OpenStudy (mertsj):

So I think it is that calories lost = calories gained It takes 1 calorie to raise 1 g of water 1 degree so to lower 85 g of water 36 degrees would require 3060 calories which must be provided by x g of water which will absorb 14x calorie. So I think x = 219 ml

OpenStudy (mertsj):

@electrokid 1 ml of water = 1 g of water at 4 degrees celsius

OpenStudy (anonymous):

@Mertsj right. You solved it logically approaching. That aint wrong But some teachers, want to see that particular equation in there; otherwise, no points. :)

OpenStudy (osuxbucks15):

Can you set it up, so i can see it in the order of that equation?

OpenStudy (osuxbucks15):

how did you get 36 degrees

OpenStudy (anonymous):

Heat lost = Heat gained \[m_{present}\times c\times (65-29)=m_{added}\times c \times (29-15)\]

OpenStudy (anonymous):

\[\cancel{\rho}\times85\times\cancel{c}\times 34=\cancel{rho}\times v\times\cancel{c}\times 14\\ v={85\times34\over14}=?{\rm ml} \]

OpenStudy (osuxbucks15):

So do i have to do anything with the heat lost= heat gained?

OpenStudy (anonymous):

thats how I got the equations

OpenStudy (osuxbucks15):

where did you get the 34

OpenStudy (anonymous):

look at the step right before that.

OpenStudy (anonymous):

found it?

OpenStudy (osuxbucks15):

65-29=36 is that it?

OpenStudy (anonymous):

oh. my bad. yea.. 36 :)

OpenStudy (osuxbucks15):

o ok so then it would be 218.6

OpenStudy (anonymous):

right. just like what @Mertsj got

OpenStudy (osuxbucks15):

ok thank you

OpenStudy (anonymous):

yw.